# PIM-malloc: A Fast and Scalable Dynamic Memory Allocator for Processing-In-Memory (PIM) Architectures

### v2

@@ 4.2. Hardware/Software Co-Designed PIM-malloc-HW/SW
Workflow. Figure 13 illustrates the workflow of PIM-malloc-HW/SW during a buddy cache hit and a miss. Initially, PIM-malloc-HW/SW utilizes the lookup_bc instruction to check whether the buddy cache stores the metadata associated with the corresponding DRAM address ( ). If the buddy cache returns a positive value, it corresponds to a buddy cache hit scenario (Figure 13(a)), whereas a negative value return corresponds to a cache miss (Figure 13(b)). When buddy cache hits, the memory allocator executes a read_bc instruction to retrieve the corresponding metadata ( ). Conversely, during a buddy cache miss, the memory allocator first retrieves the metadata from DRAM and temporarily stores it inside the on-chip scratchpad ( ), evicts the LRU entry from the buddy cache ( ), and then fills in the newly fetched metadata into the now vacant buddy cache entry using the write_bc instruction ( ). This ensures the requested metadata is readily available from the buddy cache for subsequent metadata access by the PIM core ( ).
5. Methodology
PIM core architecture | ||
Operating frequency |
Microarchitecture |
Data (WRAM) / Instruction (IRAM) scratchpad |
DRAM system | ||
DRAM bank size |
Timing Parameter |
Memory scheduling policy |
Row buffer size |
PIM-malloc configuration | ||
Per-bank DRAM heap size |
Size classes |
Buddy cache size / access latency |
@@ 5. Methodology
Characterization using a real PIM system. All the experiments conducted in Section 3 and Section 4 (with the exception of Figure 7(b), based on simulator (Hyun et al., 2024), due to the lack of profiling tools) were conducted on a real UPMEM-PIM system. This system includes an Intel Xeon Gold 5222 CPU as the host processor, equipped with two channels of DDR4-3200 and four channels of DDR-2400-based UPMEM-PIM DIMMs, containing a total of 512 PIM cores.
@@
Evaluation using cycle-level simulation. We use the open-source UPMEM-PIM cycle-level simulator, uPIMulator (Hyun et al., 2024), for all our evaluation in Section 6 that compare the performance of straw-man PIM buddy allocator, PIM-malloc-SW, and PIM-malloc-HW/SW. For PIM-malloc-HW/SW, we faithfully model the behavior of the buddy cache, the support for ISA extensions, and their HW/SW interface inside uPIMulator. Table 3 summarizes the key parameters of the PIM core architecture and PIM-malloc.
@@
Microbenchmark. In Section 6.1, we evaluate PIM-malloc’s standalone performance by comparing the average latency of the straw-man PIM buddy allocator, PIM-malloc-SW, and PIM-malloc-HW/SW. Our microbenchmark performs a series of pimMalloc() calls to measure latency while varying the number of threads and allocation sizes.
Benchmark for PIM-malloc use case. To demonstrate a practical use case of our PIM-malloc design, our evaluation in Section 6.2 focuses on a dynamic graph update application. We follow the evaluation methodology used in prior work (Awad et al., 2020; Basak et al., 2020, 2021; Pandey et al., 2021; Winter et al., 2018; Dhulipala et al., 2019) on dynamic graph updates. Due to the limited availability of open-source datasets that model dynamic graph update behaviors in real-world applications, these prior studies perform experiments using synthetic datasets. Specifically, nodes or edges of a static graph dataset are randomly sampled; the sampled portion represents the newly added graph dataset, while the unsampled portion represents the existing graph dataset. As such, we set the size ratio of the newly added dataset to the existing dataset at 1:2 and perform random sampling on the edges. We use the loc-gowalla (Cho et al., 2011) dataset, employed in the PrIM (Gómez-Luna et al., 2021b; pri, 2021) benchmark suite, as our graph dataset.
6. Evaluation
6.1. Microbenchmark based Analysis
@@
Average memory allocation latency. We compare the memory allocation latency of the straw-man PIM buddy allocator, PIM-malloc-SW, and PIM-malloc-HW/SW using the microbenchmark discussed in Section 5. In this experiment, the microbenchmark employs a single thread (Figure 14(a)) and 16 threads (Figure 14(b)) to invoke allocation requests. Each thread calls the pimMalloc() function 128 times in sequence, with each call requesting a fixed data allocation size. Generally, the straw-man PIM buddy allocator exhibits the highest latency due to its need to traverse a large tree data structure for every allocation request. For smaller allocation sizes of 32 B and 256 B, both PIM-malloc designs utilize the thread cache to achieve O(1) memory allocation latency for free memory block identification, resulting in significant performance improvements. Interestingly, PIM-malloc-HW/SW achieves faster allocations than PIM-malloc-SW even for the smaller 32 B and 256 B allocations, which technically should all be handled at the frontend thread cache. The reason for PIM-malloc-HW/SW’s advantage becomes evident in scenarios involving a thread cache miss (Figure 9(b)), where the frontend has no available free blocks, and a tree search at the backend buddy allocator is required. Under such scenarios, PIM-malloc-HW/SW can accelerate the process of retrieving the 4 KB free block from the backend buddy allocator by using its buddy cache to reduce the latency for buddy allocator’s tree traversal. For allocation requests of 4 KB, these are serviced directly by the buddy allocator for all three design points, so their performance differences primarily come from how efficiently each design’s buddy allocator is implemented. PIM-malloc-HW/SW achieves the highest performance among all three design points, thanks to its hardware-based buddy cache design, achieving an average 39% reduction in latency for 4 KB requests vs. PIM-malloc-SW. Overall, the software-only PIM-malloc-SW provides a 66 speedup compared to the straw-man PIM buddy allocator. PIM-malloc-HW/SW provides further performance improvement over PIM-malloc-SW, achieving a performance speedup.
@@
PIM-malloc-HW/SW’s sensitivity to buddy cache size. Figure 15 shows PIM-malloc-HW/SW’s speedup over PIM-malloc-SW and the corresponding buddy cache hit rate. Both the speedup and the buddy cache hit rate saturate beyond a cache size of 64 B, which can be explained by the metadata size of the buddy cache. In our buddy allocator implementation, each tree node uses 2 bits of metadata to track the allocation status of a node (tracking three states, i.e., fully allocated, partially allocated, unallocated), so a 64 B buddy cache can store metadata for up to 256 node elements (). This 64 B capacity was shown to be sufficient to capture the locality in frequently traversed tree paths, resulting in PIM-malloc-HW/SW’s speedup saturating beyond this design point. However, when the buddy cache size is reduced to less than 64 B, the shared nature of the buddy cache among concurrently running threads reduces the per-thread cache capacity. This limitation leads to significant buddy cache thrashing, resulting in frequent data transfers from DRAM and consequently degrading memory allocation performance.
6.2. Case Study: Dynamic Graph Updates
In this section, we evaluate the performance of dynamic graph updates. Figure 16(a) compares the performance of implementations based on static (CSR) and dynamic data structures (an array of linked lists (Winter et al., 2018)), where the latter is implemented using different types of dynamic memory allocators. The dynamic data structure, built with the straw-man PIM buddy allocator, exhibits lower performance compared to the static baseline, primarily due to excessive “busy-waiting”. PIM-malloc-SW alleviates this bottleneck by employing a lock-free thread cache for small allocations, thereby reducing contention and increasing computation time (“Run”). PIM-malloc-HW/SW further improves performance by using a buddy cache that accelerates metadata retrieval, allowing it to outperform the straw-man PIM buddy allocator design.
@@ 6.2. Case Study: Dynamic Graph Updates
Figure 16(b) shows the changes in memory allocation latency over time during dynamic graph updates. The straw-man PIM buddy allocator consistently exhibits the highest latency, with significant fluctuations caused by frequent inter-thread lock contention. In contrast, both PIM-malloc-SW and PIM-malloc-HW/SW maintain relatively stable and lower latencies, while PIM-malloc-HW/SW consistently outperforms PIM-malloc-SW. Occasional latency spikes occur in both designs which correspond to thread cache misses, where requests fall back to the slower buddy allocator. These results highlight the significant latency disparity between the frontend and the backend memory allocator. Notably, PIM-malloc-HW/SW effectively reduces these latency spikes compared to PIM-malloc-SW, showcasing the efficacy of our hardware-software co-design. Figure 16(c) further shows the advantages of PIM-malloc-HW/SW’s fine-grained buddy cache management and LRU replacement policy which reduce aggregate DRAM transfer size by 33% compared to PIM-malloc-SW’s software-managed metadata buffer.
6.3. Implementation Overhead
@@
We utilize CACTI 7.0 (Muralimanohar et al., 2009) with a 32 nm logic process technology to evaluate the area, power, and timing overheads of the buddy cache. The DRAM process used to implement PIM cores is approximately less dense and slower than the logic process (Devaux, 2019), so we scale the overhead of our buddy cache accordingly. The buddy cache is implemented as a 16-entry fully associative cache, with each entry comprising a 1-bit valid flag, a 4-byte tag, and a 4-byte cache block. Moreover, an LRU-based replacement policy is employed. Together, these components results in a minimal area overhead of 0.019 mm2, a power overhead of 5 mW, and an access latency of less than one PIM core logic cycle.
7. Related Work
@@
System software support for PIM systems. Several recent studies have explored system software support for PIM architectures (Wang et al., 2023a; Oliveira et al., 2024a; Lee et al., 2024b, a; Chen et al., 2023; Noh et al., 2024; Oliveira et al., 2023; Gómez-Luna et al., 2023c; Hajinazar et al., 2021; Zhao et al., 2024; Kim et al., 2023a; Giannoula et al., 2024, 2021; Oliveira et al., 2024b). SimplePIM (Chen et al., 2023) presents a software framework that enhances UPMEM-PIM system programmability by introducing abstractions akin to those found in distributed frameworks like MapReduce. UM-PIM (Zhao et al., 2024) proposes a virtual-to-physical address translation mechanism, enabling CPU and PIM cores to share a unified physical address space. Syncron (Giannoula et al., 2021) provides a hardware/software co-design to support synchronization for PIM. PUMA (Oliveira et al., 2024a) introduces a subarray-level PIM (Seshadri et al., 2015, 2016, 2013; Hajinazar et al., 2021; Seshadri et al., 2017; Chang et al., 2016; Ferreira et al., 2022; Yüksel et al., 2024; Gao et al., 2019) aware memory allocator, which supports the memory object to be placed within the specific DRAM subarrays. To the best of our knowledge, this work is the first to systematically address the unique challenges of dynamic memory allocation in commodity general-purpose PIM systems.
@@
Dynamic memory management for parallel computing. Numerous studies have been conducted to improve the performance of parallel memory (de)allocation across different threads in multi-core CPUs (Evans, 2006; Berger et al., 2000; Ghemawat and Menage, 2009; Bonwick and Adams, 2001) and GPUs (Winter et al., 2021; Pham et al., 2022; McCoy and Pandey, 2024; Gelado and Garland, 2019; Huang et al., 2010). Hoard (Berger et al., 2000) proposes a scalable allocator aimed at efficiently handling multi-threaded CPU workloads by introducing a per-thread heap (which each thread can uniquely access). XMalloc (Huang et al., 2010) presents a scalable, lock-free dynamic memory allocator designed for GPUs, addressing the scalability limitations of traditional Compare-And-Swap based allocators by leveraging SIMD-aware allocation strategies and hierarchical caching mechanisms. While prior work shares certain design principles with PIM-malloc-SW, PIM-malloc uniquely co-designs hardware/software for heterogeneous (host processor and PIM) systems, exploring optimal metadata placement and a lightweight accelerator which distinguishes itself from existing approaches.
8. Conclusion

### v3

@@ IV-B Hardware/Software Co-Designed PIM-malloc-HW/SW
Workflow. Figure 14 illustrates the workflow of PIM-malloc-HW/SW. Initially, PIM-malloc-HW/SW utilizes the lookup_bc instruction to check whether the buddy cache stores the metadata associated with the corresponding DRAM address ( ). If the buddy cache returns a positive value, it corresponds to a buddy cache hit scenario (Figure 14(a)), whereas a negative value return corresponds to a cache miss (Figure 14(b)). When buddy cache hits, the memory allocator executes a read_bc instruction to retrieve the corresponding metadata ( ). Conversely, during a buddy cache miss, the memory allocator first retrieves the metadata from DRAM and temporarily stores it inside the on-chip scratchpad ( ), evicts the LRU entry from the buddy cache ( ), and then fills in the newly fetched metadata into the now vacant buddy cache entry using the write_bc instruction ( ). This ensures that the requested metadata is readily available from the buddy cache for subsequent metadata access by the PIM core ( ).
V Methodology
@@
Characterization using a real PIM system. All the experiments conducted in Section III and Section IV (with the exception of Figure 8(b), based on simulator [64], due to the lack of profiling tools) were conducted on a real UPMEM-PIM system. This system includes an Intel Xeon Gold 5222 CPU as the host processor, equipped with two channels of DDR4-3200 and four channels of DDR-2400-based UPMEM-PIM DIMMs, containing a total of 512 PIM cores.
@@
Evaluation using cycle-level simulation. We use the open-source UPMEM-PIM cycle-level simulator, uPIMulator [64], for all our evaluation in Section VI that compare the performance of straw-man PIM buddy allocator, PIM-malloc-SW, and PIM-malloc-HW/SW. For PIM-malloc-HW/SW, we faithfully model the behavior of the buddy cache, the support for ISA extensions, and their HW/SW interface inside uPIMulator. We adopt the default simulation configuration provided by uPIMulator [64] and configure PIM-malloc as follows: a 32 MB heap size; allocation size classes defined by powers of two (with a minimum of 16 B and a maximum of 2 KB, total of 8 size classes); a buddy cache size of 64 B; and a buddy cache access latency of 1 PIM core cycle.
@@
Microbenchmark. In Section VI-A, we evaluate PIM-malloc’s standalone performance by comparing the average latency of the straw-man PIM buddy allocator, PIM-malloc-SW, and PIM-malloc-HW/SW. Our microbenchmark performs a series of pimMalloc() calls to measure latency while varying the number of threads, allocation size, etc.
Benchmark for PIM-malloc use case. To demonstrate a practical use case of our PIM-malloc design, we evaluate (1) dynamic graph update and (2) attention layer in LLM inference. For the dynamic graph update, we follow the evaluation methodology used in prior work [5, 7, 8, 109, 134, 30]. Due to the limited availability of open-source datasets that model dynamic graph update behaviors in real-world applications, these prior studies perform experiments using synthetic datasets. Specifically, nodes or edges of a static graph dataset are randomly sampled; the sampled portion represents the newly added graph dataset, while the unsampled portion represents the existing graph dataset. As such, we set the size ratio of the newly added dataset to the existing dataset at 1:2 and perform random sampling on the edges. We use the loc-gowalla [23] dataset, employed in the PrIM [52] benchmark suite, as our graph dataset. It is important to note that the performance of dynamic graph updates varies by data structure. Therefore, we evaluate two representative structures from prior work: an array of linked lists [134] and a variable-sized array [16]. The former uses a constant allocation size (we assume 256 B) because its edge-storing elements are fixed-size arrays. The latter manages all edges for a node in a single power-of-two-sized array, resulting in allocation sizes that vary from 64 B to 32 KB in our dataset. For the attention layer, we adopt the Llama‑2 7B [129] model configuration to calculate the KV cache size and extend the PrIM [52] GEMV benchmark to execute the attention computation on PIM. To handle the dynamic growth of the KV cache, our PIM kernel is designed to allocate a new 512 B memory block whenever the existing space is exhausted. We evaluate LLM inference performance by generating a trace using simulation results from uPIMulator [64]. This trace represents a workload of 100 total requests arriving at a rate of 10 requests per second, each with an input prompt of 128 tokens and an output of 256 tokens. This trace is then processed by LLMServingSim [24], a system-level simulator for LLM inference serving.
VI Evaluation
VI-A Microbenchmark based Analysis
@@
Average memory allocation latency. We compare memory allocation latency across various allocators using the microbenchmark described in Section V. The microbenchmark uses either one (Figure 15(a)) or 16 threads (Figure 15(b)), each making 128 allocation requests of a fixed size. The straw-man allocator has the highest latency, since it needs to traverse a large tree structure for every request. For smaller allocations (32 B and 256 B), both PIM-malloc designs rely on a thread cache to achieve O(1) latency, yielding significant speedups. PIM-malloc-HW/SW provides further speedup over PIM-malloc-SW for 32 B and 256 B allocations because, in cases where the thread cache is empty, its hardware-based buddy cache accelerates tree traversal. For allocation requests of 4 KB, these are serviced directly by the buddy allocator for all three design points, so their performance differences primarily come from how efficiently each design’s buddy allocator is implemented. PIM-malloc-HW/SW’s hardware-based cache reduces 4 KB allocation latency by 39% compared to PIM-malloc-SW. Overall, PIM-malloc-SW achieves a 66 speedup over the straw-man allocator, while PIM-malloc-HW/SW outperforms PIM-malloc-SW by 31%.
@@
PIM-malloc-HW/SW’s sensitivity to buddy cache size. Figure 16 shows PIM-malloc-HW/SW’s speedup over PIM-malloc-SW and the corresponding buddy cache hit rate. Both the speedup and the buddy cache hit rate saturate beyond a cache size of 64 B, which can be explained by the metadata size of the buddy cache. In our buddy allocator implementation, each tree node uses 2 bits of metadata to track the allocation status of a node (tracking three states, i.e., fully allocated, partially allocated, unallocated), so a 64 B buddy cache can store metadata for up to 256 node elements (). 64 B capacity was shown to be sufficient to capture the locality in frequently traversed tree paths, resulting in PIM-malloc-HW/SW’s speedup saturating beyond this design point.
Comparison of metadata management in PIM-malloc-SW and PIM-malloc-HW/SW. The buddy allocator traverses a buddy tree to access metadata, creating a non-sequential access pattern as it probes physically non-adjacent entries. This characteristic renders the coarse-grained caching scheme of PIM-malloc-SW inefficient. This scheme fetches large, often unrelated metadata blocks, which wastes MRAM-to-WRAM bandwidth. Consequently, it suffers from a low buffer hit rate and a high volume of data transfers caused by frequent block replacements. In contrast, the fine-grained approach of PIM-malloc-HW/SW effectively leverages the buddy allocator’s temporal locality through a hardware-based LRU policy. It minimizes unnecessary data transfers by retaining the metadata of frequently revisited parent nodes from upper tree levels in its buddy cache. The results from a 4 KB allocation microbenchmark clearly illustrate this difference. PIM-malloc-HW/SW achieved a 99% hit rate, transferring on average only 2-bytes of metadata per pimMalloc() request, whereas PIM-malloc-SW’s 73% hit rate resulted in 2 KB of data transfer per pimMalloc() request. However, this performance gap narrows under certain conditions, such as with large allocation sizes (e.g., 256 KB) and a low number of requests. This is because the shallow search depth required for large allocations increases metadata reuse within a single coarse-grained block, and the low request frequency makes buffer replacements infrequent.
VI-B Case Study #1: Dynamic Graph Updates
@@
In this section, we evaluate the performance of dynamic graph updates. Figure 17(a) compares the performance of implementations based on static (CSR) and dynamic data structures (array of linked lists and variable-sized arrays), where the latter is implemented using different types of dynamic memory allocators. The dynamic data structure, built with the straw-man PIM buddy allocator, exhibits lower performance compared to the static baseline, primarily due to excessive “busy-waiting”. PIM-malloc-SW alleviates this bottleneck by employing a lock-free thread cache for small allocations, thereby reducing contention and increasing the computation time ratio (“Run”). PIM-malloc-HW/SW further improves performance by using a buddy cache that accelerates metadata retrieval, allowing it to outperform the straw-man PIM buddy allocator design. Consequently, PIM-malloc-HW/SW demonstrates a 7.1 and 32 speedup over the static baseline for the array of linked lists and the variable-sized array implementations, respectively. Figure 17(b) presents the allocation latency distribution during dynamic graph updates. The straw-man allocator suffers from extremely high average latencies, on the order of hundreds of milliseconds, which are omitted from the figure for readability. By comparison, both PIM-malloc-SW and PIM-malloc-HW/SW drastically reduce this latency, underscoring that minimizing allocation overhead is critical for achieving high workload performance. Figure 17(c) shows the changes in memory allocation latency over time during dynamic graph updates. The straw-man PIM buddy allocator consistently exhibits the highest latency, with significant fluctuations caused by frequent inter-thread lock contention. In contrast, both PIM-malloc-SW and PIM-malloc-HW/SW maintain relatively stable and lower latencies, while PIM-malloc-HW/SW consistently outperforms PIM-malloc-SW. Occasional latency spikes occur in both designs which correspond to thread cache misses, where requests fall back to the slower buddy allocator. These results highlight the significant latency disparity between the frontend and the backend memory allocator. Figure 17(d) further shows the advantages of PIM-malloc-HW/SW’s fine-grained buddy cache management and LRU replacement policy which reduce aggregate DRAM transfer size by 30% compared to PIM-malloc-SW’s software-managed metadata buffer.
VI-C Case Study #2: Attention Layer in LLM Inference
Two key metrics measure the performance of LLM serving: (1) Throughput, the number of tokens processed per second, and (2) Time-Per-Output-Token (TPOT), the latency between generating consecutive tokens. As explained in Section III-A, managing the KV cache for LLM inference with dynamic allocation enables more efficient memory usage. This allows for more concurrent requests, resulting in higher throughput compared to static allocation (Figure 18). Furthermore, reducing the dynamic allocator’s latency frees up PIM’s computational resources for the attention operation, further improving throughput. Among the dynamic methods, PIM-malloc-HW/SW delivers the highest throughput (1.7 improvement over static allocation). However, this approach presents a trade-off: dynamic memory allocation introduces overhead that increases the attention layer’s latency, which in turn increases TPOT. This trend is observed consistently across all TPOT percentiles. The straw-man method, for instance, exhibits the highest TPOT among the four schemes due to its high allocation latency. In contrast, PIM-malloc-SW effectively achieves a lower TPOT than the straw-man approach via its thread cache (frontend) design. Building on this, PIM-malloc-HW/SW leverages its buddy cache to improve both the 95th and 99th percentile TPOT by another 8% over PIM-malloc-SW, resulting in TPOTs that are 21% and 24% higher than the static allocation baseline, respectively.
VI-D Fragmentation Analysis
PIM-malloc mitigates fragmentation by employing dedicated memory pools for each size class. However, to reduce initial allocation latency, the current implementation includes an optimization that pre-allocates all thread caches with free 4 KB memory blocks during the initAllocator() call, as discussed in Section IV-A. While this pre-population strategy can enhance performance, it may also lead to unnecessary memory fragmentation, depending on the workload’s allocation patterns. For example, if threads on a PIM core only allocate memory blocks of a specific size, the pre-allocated memory in other size classes remains entirely unused.
Table III quantifies memory fragmentation (/), measured as the ratio between “total allocated memory by allocator ()” and “total memory requested by program ()”, as defined in [9], e.g., a fragmentation value higher than 1.0 means that memory is reserved but not effectively used by the program. The dynamic graph update (array of linked lists) workload, which uses only a single size class, exhibits a high fragmentation of 1.95 due to its unused pre-allocated memory blocks. In contrast, fragmentation decreases to 1.72 for the dynamic graph update (variable sized array), which utilizes a variety of block sizes. Interestingly, the LLM attention workload shows even lower fragmentation at 1.66. Although it also uses a single allocation size, its total allocated volume is large enough to amortize the unused pre-allocated blocks. To address this problem, we evaluate a version of our allocator with pre-allocation disabled (PIM-malloc-lazy in Table III). This “lazy” approach, which allocates memory on-demand rather than pre-filling thread caches, shows a significant reduction in fragmentation across all workloads. These findings highlight a clear trade-off between initial allocation performance and overall memory efficiency. We leave the optimization of this balance, potentially through compiler-based analysis or a user-configurable interface, as future work.
| Workload | PIM-malloc (as-is) | PIM-malloc-lazy |
| Dynamic graph update | 1.95 | 1.21 |
| (Array of linked list) | ||
| Dynamic graph update | 1.72 | 1.49 |
| (Variable sized array) | ||
| LLM attention | 1.66 | 1 |
VI-E Metadata Storage Overhead of PIM-malloc
According to our analysis, the metadata storage overhead of PIM-malloc is minimal. Its hierarchical structure reduces the buddy tree depth to 13 compared to a straw-man design, shrinking the metadata to just 4 KB per DRAM bank. PIM-malloc also uses a bitmap for the thread cache’s metadata, making its storage cost negligible. Across our workloads, the maximum overhead is 5.1 KB for dynamic graph update (array of linked lists), 5 KB for dynamic graph update (variable-sized array), and 5.2 KB for LLM attention per request.
VI-F Implementation Overhead
@@
We use CACTI 7.0 [96] with a 32 nm logic process technology to evaluate the area, power, and timing overheads of the buddy cache. The DRAM process used to implement PIM cores is approximately less dense and slower than the logic process [29], so we scale the overhead of our buddy cache accordingly. Our evaluation indicates that the buddy cache adds only 0.019 mm2, a power overhead of 5 mW, and an access latency of less than one PIM core logic cycle.
VII Discussion
Applicability of PIM-malloc to future PIM architectures. We consider two potential architectural evolution paths inspired by recent work by Hyun et al. [64], which discussed future PIM capable of exploiting higher parallelism and locality. First, consider a future PIM device with enhanced processing capabilities enabled by DRAM process optimizations. Although such advancements would reduce the absolute latency of pimMalloc(), they would also proportionally accelerate the overall workload, implying that the relative overhead of memory allocation would likely remain a key performance bottleneck. Second, in cache-enabled PIM architectures, even with a general-purpose data cache, a specialized metadata cache, such as our proposed buddy cache, remains necessary. This need stems from a fundamental granularity mismatch: general-purpose caches operate on coarse-grained cache lines (e.g., 64 bytes), which is inefficient for managing the fine-grained metadata used by a buddy allocator. Our buddy cache addresses this by handling metadata at a much finer granularity (e.g., 8 bytes), complementing the general-purpose cache and addressing the memory allocation bottleneck.
VIII Related Work
@@
System software support for PIM systems. Several recent studies have explored system software support for PIM architectures [132, 105, 84, 83, 20, 102, 106, 51, 57, 141, 72, 44, 43, 107]. SimplePIM [20] presents a software framework that enhances UPMEM-PIM system programmability by introducing abstractions akin to those found in distributed frameworks. Syncron [43] provides a hardware/software co-design to support synchronization for PIM. PUMA [105] introduces a subarray-level PIM [118, 120, 119, 57, 121, 19, 34, 138, 36] aware memory allocator, which supports the memory object to be placed within the specific DRAM subarrays. To the best of our knowledge, this work is the first to systematically address the unique challenges of dynamic memory allocation in general-purpose PIM systems.
@@
Dynamic memory management for parallel computing. Numerous studies have been conducted to improve the performance of parallel memory (de)allocation across different threads in multi-core CPUs [32, 9, 40, 11] and GPUs [135, 113, 95, 39, 63]. Hoard [9] proposes a scalable allocator aimed at efficiently handling multi-threaded CPU workloads by introducing a per-thread heap (which each thread can uniquely access). XMalloc [63] presents a scalable, lock-free dynamic memory allocator designed for GPUs, addressing the scalability limitations of traditional Compare-And-Swap based allocators by leveraging SIMD-aware allocation strategies and hierarchical caching mechanisms. While prior work shares certain design principles with PIM-malloc-SW, PIM-malloc uniquely co-designs hardware/software for heterogeneous (host processor and PIM) systems, exploring optimal metadata placement and a lightweight accelerator which distinguish itself from existing approaches.
IX Conclusion